# The Power of Iterative Filtering for Supervised Learning with (Heavy) Contamination

### v1

@@ Abstract
    We show that any function class that can be approximated by low-degree polynomials with respect to a hypercontractive distribution can be efficiently learned under bounded contamination (also known as nasty noise). This is a surprising resolution to a longstanding gap between the complexity of agnostic learning and learning with contamination, as it was widely believed that low-degree approximators only implied tolerance to label noise.
@@ 1 Introduction
[…]
@@
In Theorem 4.2, we give a positive answer to this question, thereby resolving a longstanding gap between the complexity of learning with bounded contamination and with adversarial label noise. This is particularly important as it implies exponential improvements for BC-learning of fundamental concept classes like intersections of halfspaces, monotone functions, and convex sets (see Table 2).
@@
    Then, an adversary receives $\bar{S}_{\mathrm{cln}}$ and adds $M-m$ arbitrary labeled examples to form $\bar{S}_{\mathrm{inp}}$.
The heavy contamination model only allows the adversary to add points, since removing an arbitrary fraction of the clean samples would correspond to learning with truncation [23, 24, 80], which is beyond the scope of this work (see also Remark 4.6). Another difference between Definition 1.1 and Definition 1.3 is that in the HC model, clean labels need not be realized by some function in the given concept class. We instead consider the following quantity
@@
Proposition 1.5 (Informal, see Propositions B.3 and B.2).
@@
Naturally, one might wonder whether the existence of approximating polynomials is sufficient for HC-learning. In Theorem 5.2 we give a negative answer to this question, by providing a lower bound on the sample complexity of HC-learning of monotone functions, which admit low-degree approximators. A recent line of works has used the stronger notion of sandwiching approximators to provide efficient algorithms for various challenging learning tasks [48, 78, 54, 18, 79]. Here, we expand on this paradigm and show that the existence of low-degree sandwiching polynomials implies efficient HC-learning as well (Theorem 4.4).
@@ 1.1 Our Results
In Table 1, we present an overview of the upper bounds we obtain as applications of our main theorems on learning with contamination (Theorems 4.2 and 4.4). For comparison, we also provide the corresponding results on learning with label noise from prior work. For the case of bounded contamination, and for constant error and confidence parameters ($\epsilon,\delta$), our results match the best known bounds for learning with label noise, since in both cases, the existence of low-degree polynomial approximators is sufficient. See Tables 2 and 3 for further details.
@@
Table 1: Upper and lower bounds on the time complexity of learning in different noise models up to excess error $\epsilon=0.1$ and failure probability $\delta=0.01$. See Appendix D for more details on the complexity of learning with adversarial label noise.
All of our results work in the distribution-specific setting. Distributional assumptions are unavoidable, as there is strong evidence that distribution-free learning of even the simplest classes (e.g., linear classifiers) in the presence of noise is computationally hard [44, 22]. Although we present our results for specific standard target marginal distributions, our results hold for any hypercontractive (Definition 3.1) marginal $\mathcal{D}^{*}$ that can be sampled efficiently, as long as the degree of approximation (resp. sandwiching for HC-learning) of the learned class is low under $\mathcal{D}^{*}$. Hypercontractive distributions are an extremely wide class of probability distributions, which includes Gaussians, all log-concave distributions over ${\mathbb{R}}^{d}$, as well as product distributions over {±$\{\pm 1\}^{d}$. See Appendix C for an overview of relevant results in approximation theory.
On the lower bound side, we show that learning monotone functions with heavy contamination requires exponentially many samples, thereby separating bounded and heavy contamination. For a more thorough discussion on lower bounds for learning with contamination, see Section 5.222We note that although learning with bounded contamination is clearly more challenging than learning with adversarial label noise, many of the known lower bounds for agnostic learning do not transfer directly to BC-learning as formalized in Definition 1.1 for technical reasons. See Section 5 for a way to circumvent this.
@@
A recent line of works in learning theory has focused on providing learning algorithms that can verify their distributional assumptions [83, 48, 50, 33, 51, 87, 54]. These algorithms are allowed to either accept and output a classifier with certified optimal performance, or detect a violation of their target distributional assumptions and reject. Here, we provide improved results for a tolerant version of this problem where the algorithm has to accept even if the input distribution is close to the target distribution. Tolerant testable learning was first studied in [54] (see Definition 4.7).
In Theorem 4.8, we show that the existence of low-degree sandwiching polynomials implies efficient tolerant testable learning. Previously, [54] required the existence of low-degree $\mathcal{L}_{2}$-sandwiching polynomials, which is a stronger assumption. Moreover, all of the previous results in testable learning—even the non-tolerant variants—required that the sandwiching polynomials have bounded coefficients. Here, we do not impose such a requirement and obtain the first testable learning results for functions of halfspaces with respect to any fixed log-concave distribution. All prior work either gave worse error guarantees [51, 50], or required target marginals with strictly sub-exponential tails [48]. See Section 4.3 for more details.
@@ 1.2 Our Techniques
All of our results use the same iterative polynomial filtering algorithm of Theorem 3.2 with appropriate hyperparameter choices. The algorithm receives a set $S_{\mathrm{inp}}$ of data points and filters it, outputting a subset $S_{\mathrm{filt}}$ of $S_{\mathrm{inp}}$ that satisfies the following two conditions. First, any low-degree polynomial $p$ whose absolute expectation $\operatorname*{\mathbb{E}}[|p(\mathbf{x})|]$ over the pre-specified target distribution $\mathcal{D}^{*}$ is small, will also have bounded average over $S_{\mathrm{filt}}$, i.e. $\frac{1}{|S_{\mathrm{filt}}|}\sum_{\mathbf{x}\in S_{\mathrm{filt}}}p(\mathbf{x})$ is small. Second, if $S_{\mathrm{inp}}$ contains a set of points $S$ that were initially generated independently by $\mathcal{D}^{*}$, then only a limited number of points from $S$ can be removed by the filtering (the allowed number of removed points is controlled by an appropriate hyperparameter).
@@
Hence, the first phase of our algorithm aims to filter out such bad input datapoints. The following basic observation is key to our approach: if $p^{*}$ is an $\epsilon^{4}$-approximator in $\mathcal{L}_{2}$ norm for a {±$\{\pm 1\}$-valued function $f^{*}$, then the average $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}^{*}}[|(p^{*}(\mathbf{x}))^{2}-1|]$ is at most $O(\epsilon^{2})$. This observation, together with our filtering guarantee, ensures that the average $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[(p^{*}(\mathbf{x}))^{2}-1]$ is likewise bounded by $O(\epsilon)$ while removing almost exclusively outliers.
@@ 1.3 Related Work
Our Definition 1.3 is inspired by semi-random models, which lie between the average and worst case settings [13, 45]. In particular, it resembles an instantiation of the semi-random model framework known as a monotone adversary, which breaks a statistical assumption used by a learning algorithm (e.g., i.i.d. draws from a known distribution) by providing additional data. Our approach is inspired by recent algorithms for supervised regression problems, e.g., solving linear systems, sparse recovery, or matrix completion [16, 70, 59, 71], that are tolerant to monotone adversaries. These algorithms also use reweightings that come with certificates of success. However, a major qualitative difference between the aforementioned works and ours is that our Definition 1.3 does not require that the adversary uses labels consistent with a “clean hypothesis.” Instead, our algorithms can tolerate label noise (alongside covariate noise) and achieve the information-theoretically optimal clean error under such a model (Proposition 1.5).
@@
Learning with heavy contamination can be thought of as a search version of testable learning. More specifically, in testable learning the goal is to decide whether the input dataset is structured enough so that a near-optimal hypothesis can be found efficiently, while in HC-learning the goal is to find a subset of the input that is structured enough. An analogous connection was observed in [54] between TDS learning [78] and PQ learning [49] in the context of learning under distribution shift. There, the goal was to either decide whether the (unlabeled) test examples come from a distribution that is similar to the one the learner has trained on (TDS), or find a subset of the test examples where the learner is confident in its predictions (PQ). See Remark 4.5.
@@ 2 Notation
We consider a $d$-dimensional feature space $\mathcal{X}$ which will either be ${\mathbb{R}}^{d}$ or the hypercube {±$\{\pm 1\}^{d}$. A polynomial $p$ over ${\mathbb{R}}^{d}$ is of the form $p(\mathbf{x})=\sum_{\alpha\subseteq{\mathbb{N}}^{d}}c_{p}(\alpha)\,\mathbf{x}^{\alpha}$, where $\mathbf{x}^{\alpha}=\prod_{i\in[d]}x_{i}^{\alpha_{i}}$ is a monomial of degree $\|\alpha\|_{1}$. The degree of $p$ is equal to the maximum $\|\alpha\|_{1}$ such that $c_{p}(\alpha)\neq 0$. Over {±$\{\pm 1\}^{d}$ we use the multilinear expansion $p(\mathbf{x})=\sum_{\mathcal{I}\subseteq[d]}c_{p}(\mathcal{I})\mathbf{x}^{\mathcal{I}}$, where $\mathbf{x}^{\mathcal{I}}=\prod_{i\in\mathcal{I}}x_{i}$. In both cases, we call $\mathbf{c}_{p}$ the vector of coefficients. We consider sets of points $S$ to contain examples that are separate instances of the correponding elements of $\mathcal{X}$. We denote with $\bar{S}$ the corresponding labeled set of examples. We use $\mathbf{x}\sim S$ to say that $\mathbf{x}$ is drawn uniformly from $S$. For a labeled distribution $\bar{\mathcal{D}}$ over $\mathcal{X}\times\{\pm 1\}$, $\mathcal{D}$ is the marginal on $\mathcal{X}$. In the following the distribution-specific algorithms have sample access to the unlabeled target distribution $\mathcal{D}^{*}$. We use $\mathcal{N}_{d}=\mathcal{N}(0,\mathbf{I}_{d})$ to denote the standard $d$-dimensional Gaussian and $\mathrm{Unif}_{d}=\mathrm{Unif}\{\pm 1\}^{d}$ for the uniform over the $d$-dimensional hypercube.
@@ 3 Iterative Polynomial Filtering
We say that a distribution $\mathcal{D}$ over $\mathcal{X}$ is ${A}$-hypercontractive with respect to polynomials for some ${A}\geq 1$ if for any polynomial $p$ over $\mathcal{X}$ and any $t\geq 2$ we have
@@
Let $\mathcal{D}^{*}$ be a ${A}$-hypercontractive distribution over a $d$-dimensional space $\mathcal{X}$. Consider parameters $\epsilon,\delta\in(0,1)$ and $R,\ell,m\geq 1$, and let $S_{\mathrm{inp}}$ be an arbitrary set of examples in $\mathcal{X}$ and $S_{\mathrm{ref}}$ a set of $m_{\mathrm{ref}}$ i.i.d. examples from $\mathcal{D}^{*}$. For a sufficiently large universal constant $C\geq 1$, if $m_{\mathrm{ref}}\geq\frac{R^{2}(CAd)^{2\ell}}{\epsilon^{3}}(\log\frac{1}{\delta})^{4\ell+1}$, then Algorithm 1 on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m,\ell,R,\epsilon)$ runs in time $\operatorname{poly}(|S_{\mathrm{inp}}|,m_{\mathrm{ref}},(d+1)^{\ell})$ and outputs $S_{\mathrm{filt}}\subseteq S_{\mathrm{inp}}$ such that the following hold:
@@
1 Let $\beta\leftarrow 2(2A)^{2\ell}$; $\gamma\leftarrow\frac{\epsilon}{2R}$; $B\leftarrow 4(d+1)^{\frac{\ell}{2}}(\frac{\beta}{\epsilon})^{\frac{1}{2}}$; $\Delta\leftarrow\frac{\epsilon}{2B}$;
@@
We now give the proof of our main technical tool, Theorem 3.2. The proof follows the approach of [79], but has a number of technical differences. In particular, the algorithm of [79] only gives guarantees for non-negative polynomials and works with respect to the uniform distribution on the hypercube. Here, we preserve the expectation of any polynomial, with respect to any target hypercontractive distribution that can be sampled efficiently.
Proof of Theorem 3.2.
We first observe that the family $\mathcal{P}$ of polynomials $p$ of degree at most $\ell$ for which $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p(\mathbf{x})|]\leq\gamma$ and $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[(p(\mathbf{x}))^{2}]\leq\beta$ can be described by $O(m_{\mathrm{ref}})$ linear constraints plus one convex quadratic constraint over the coefficient vectors whose dimension is $(d+1)^{\ell}$. Therefore, lines 1 and 5 can be implemented as convex programs in time $\operatorname{poly}(|S_{\mathrm{inp}}|,m_{\mathrm{ref}},(d+1)^{\ell})$. Lines 1 and 1 can be implemented with a single pass of the input points.
@@
[…]
@@
We will show that with high probability over the clean and reference datasets, the number of removed points from $S_{\mathrm{cln}}$ is small. First, we account for the initial filtering of line 1. In this step, we remove points $\mathbf{x}$ from $S_{\mathrm{inp}}$ that give large absolute values to polynomials in $\mathcal{P}$, i.e., polynomials who, in particular, have bounded second norms over $S_{\mathrm{ref}}$. Due to Lemma A.2 and since $m_{\mathrm{ref}}\geq(CAd)^{2\ell}(\log\frac{1}{\delta})^{4\ell+1}$, the probability that some $\mathbf{x}$ drawn from $\mathcal{D}$ gives $|p(\mathbf{x})|>B$ for some $p\in\mathcal{P}$ is at most $\epsilon/4$. Therefore, the total number of points removed from $S_{\mathrm{cln}}$ in this step follows the binomial distribution with $m$ number of trials and probability of success at most $\epsilon/4$. By a standard Chernoff bound, and since $m\geq\frac{C}{\epsilon}\log(1/\delta)$, with probability at least $21-\delta/$ 2 we have
@@
Since $m,m_{\mathrm{ref}}\geq C^{\prime}\frac{(d+1)^{\ell}+\log(1/\delta)}{(\Delta/R)^{2}}$ for some sufficiently large universal constant $C^{\prime}\geq 1$ and the VC dimension of the class of polynomial threshold functions of degree $\ell$ is at most $(d+1)^{\ell}$, we have that with probability at least $21-\delta/$ 2 over $S_{\mathrm{cln}}$, $S_{\mathrm{ref}}$ the following holds for all polynomials $p$ of degree at most $\ell$ and all $\tau\geq 0$
@@
since $S_{\mathrm{cln}}$ and $S_{\mathrm{ref}}$ are both i.i.d. samples from the distribution $\mathcal{D}^{*}$ and $\operatorname{\mathbbm{1}}\{|p(\mathbf{x})|>\tau\}$ is equal to the sum of two $-\ell$ polynomial threshold functions, $\operatorname{\mathbbm{1}}\{p(\mathbf{x})>\tau\}+\operatorname{\mathbbm{1}}\{p(\mathbf{x})<-\tau\}$. Therefore:
@@
So far, we have proven the first part of Theorem 3.2. We will now prove the second part, by showing the following claim.

### v2

@@ Abstract
    We show that any function class that can be approximated by low-degree polynomials with respect to a hypercontractive distribution can be efficiently learned under bounded contamination (also known as nasty noise). This is a surprising resolution to a longstanding gap between the complexity of agnostic learning and learning with contamination, as it was widely believed that low-degree approximators only implied tolerance to label noise. In particular, it implies the first efficient algorithm for learning halfspaces with $\eta$-bounded contamination up to error $2\eta+\epsilon$ with respect to the Gaussian distribution.
@@ 1 Introduction
Let $\mathcal{D}$ be some distribution over $\mathcal{X}$, $\eta\in(0,1)$ and $f:\mathcal{X}\to\{\pm 1\}$. We say that a set of samples $\bar{S}_{\mathrm{inp}}$ is generated by $(\mathcal{D},f)$ with bounded contamination of rate $\eta$ if it is generated as follows for some $M\geq 1$. [^1] We use the notation $\bar{S}$ to denote a labeled dataset and distinguish it from its unlabeled counterpart $S$.
@@
In Theorem˜4.2, we give a positive answer to this question, thereby resolving a longstanding gap between the complexity of learning with bounded contamination and with adversarial label noise. This is particularly important as it implies exponential improvements for BC-learning of fundamental classes like intersections of halfspaces, monotone functions, and convex sets (see Table˜2).
Moreover, our result provides the first algorithm for learning halfspaces with $\eta$-bounded contamination up to error $2\eta+\epsilon$ with respect to the Gaussian distribution. [^2] The work of [76] achieves a similar information-theoretically optimal guarantee under the uniform distribution on the hypercube. Prior results for learning halfspaces with bounded contamination incurred strictly worse error guarantees, differing by at least a constant multiplicative factor.
@@
    Then, an adversary receives $\bar{S}_{\mathrm{cln}}$ and adds MmM-m arbitrary labeled examples to form $\bar{S}_{\mathrm{inp}}$.
The heavy contamination model only allows the adversary to add points, since removing an arbitrary fraction of the clean samples would correspond to learning with truncation [24, 25, 61], which is beyond the scope of this work (see also Remark˜4.6). Another difference between Definition˜1.1 and Definition˜1.3 is that in the HC model, clean labels need not be realized by some function in the given concept class. We instead consider the following quantity
@@
Proposition 1.5 (Informal, see Propositions˜B.3 and B.2).
@@
Naturally, one might wonder whether the existence of approximating polynomials is sufficient for HC-learning. In Theorem˜5.2 we give a negative answer to this question, by providing a lower bound on the sample complexity of HC-learning of monotone functions, which admit low-degree approximators. A recent line of works has used the stronger notion of sandwiching approximators to provide efficient algorithms for various challenging learning tasks [51, 78, 49, 16, 76]. Here, we expand on this paradigm and show that the existence of low-degree sandwiching polynomials implies efficient HC-learning as well (Theorem˜4.4).
@@ 1.1 Our Results
In Table˜1, we present an overview of the upper bounds we obtain as applications of our main theorems on learning with contamination (Theorems˜4.2 and 4.4). For comparison, we also provide the corresponding results on learning with label noise from prior work. For the case of bounded contamination, and for constant error and confidence parameters ($\epsilon,\delta$), our results match the best known bounds for learning with label noise, since in both cases, the existence of low-degree polynomial approximators is sufficient. See Tables˜2 and 3 for further details.
@@
Table 1: Upper and lower bounds on the time complexity of learning in different noise models up to excess error $\epsilon=0.1$ and failure probability $\delta=0.01$. See Appendix˜D for more details on the complexity of learning with adversarial label noise.
All of our results work in the distribution-specific setting. Distributional assumptions are unavoidable, as there is strong evidence that distribution-free learning of even the simplest classes (e.g., linear classifiers) in the presence of noise is computationally hard [44, 23]. Although we present our results for specific standard target marginal distributions, our results hold for any hypercontractive (Definition˜3.1) marginal $\mathcal{D}^{*}$ that can be sampled efficiently, as long as the degree of approximation (resp. sandwiching for HC-learning) of the learned class is low under $\mathcal{D}^{*}$. Hypercontractive distributions are an extremely wide class of probability distributions, which includes Gaussians, all log-concave distributions over ${\mathbb{R}}^{d}$, as well as product distributions over $\{\pm 1\}^{d}$. See Appendix˜C for an overview of relevant results in approximation theory.
On the lower bound side, we show that learning monotone functions with heavy contamination requires exponentially many samples, thereby separating bounded and heavy contamination. For a more thorough discussion on lower bounds for learning with contamination, see Section˜5.333We note that although learning with bounded contamination is clearly more challenging than learning with adversarial label noise, many of the known lower bounds for agnostic learning do not transfer directly to BC-learning as formalized in Definition 1.1 for technical reasons. See Section 5 for a way to circumvent this.
@@
A recent line of works in learning theory has focused on providing learning algorithms that can verify their distributional assumptions [84, 51, 53, 31, 52, 86, 49]. These algorithms are allowed to either accept and output a classifier with certified optimal performance, or detect a violation of their target distributional assumptions and reject. Here, we provide improved results for a tolerant version of this problem where the algorithm has to accept even if the input distribution is close to the target distribution. Tolerant testable learning was first studied in [49] (see Definition˜4.7).
In Theorem˜4.8, we show that the existence of low-degree sandwiching polynomials implies efficient tolerant testable learning. Previously, [49] required the existence of low-degree $\mathcal{L}_{2}$-sandwiching polynomials, which is a stronger assumption. Moreover, all of the previous results in testable learning—even the non-tolerant variants—required that the sandwiching polynomials have bounded coefficients. Here, we do not impose such a requirement and obtain the first testable learning results for functions of halfspaces with respect to any fixed log-concave distribution. All prior work either gave worse error guarantees [52, 53], or required target marginals with strictly sub-exponential tails [51]. See Section˜4.3 for more details.
@@ 1.2 Our Techniques
All of our results use the same iterative polynomial filtering algorithm of Theorem˜3.2 with appropriate hyperparameter choices. The algorithm receives a set $S_{\mathrm{inp}}$ of data points and filters it, outputting a subset $S_{\mathrm{filt}}$ of $S_{\mathrm{inp}}$ that satisfies the following two conditions. First, any low-degree polynomial $p$ whose absolute expectation $\operatorname*{\mathbb{E}}[|p(\mathbf{x})|]$ over the pre-specified target distribution $\mathcal{D}^{*}$ is small, will also have bounded average over $S_{\mathrm{filt}}$, i.e. $\frac{1}{|S_{\mathrm{filt}}|}\sum_{\mathbf{x}\in S_{\mathrm{filt}}}p(\mathbf{x})$ is small. Second, if $S_{\mathrm{inp}}$ contains a set of points $S$ that were initially generated independently by $\mathcal{D}^{*}$, then only a limited number of points from $S$ can be removed by the filtering (the allowed number of removed points is controlled by an appropriate hyperparameter).
@@
Hence, the first phase of our algorithm aims to filter out such bad input datapoints. The following basic observation is key to our approach: if $p^{*}$ is an $\epsilon^{4}$-approximator in $\mathcal{L}_{2}$ norm for a $\{\pm 1\}$-valued function $f^{*}$, then the average $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}^{*}}[|(p^{*}(\mathbf{x}))^{2}-1|]$ is at most $O(\epsilon^{2})$. This observation, together with our filtering guarantee, ensures that the average $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[(p^{*}(\mathbf{x}))^{2}-1]$ is likewise bounded by $O(\epsilon)$ while removing almost exclusively outliers.
@@ 1.3 Related Work
Our Definition˜1.3 is inspired by semi-random models, which lie between the average and worst case settings [10, 43]. In particular, it resembles an instantiation of the semi-random model framework known as a monotone adversary, which breaks a statistical assumption used by a learning algorithm (e.g., i.i.d. draws from a known distribution) by providing additional data. Our approach is inspired by recent algorithms for supervised regression problems, e.g., solving linear systems, sparse recovery, or matrix completion [20, 71, 59, 70], that are tolerant to monotone adversaries. These algorithms also use reweightings that come with certificates of success. However, a major qualitative difference between the aforementioned works and ours is that our Definition˜1.3 does not require that the adversary uses labels consistent with a “clean hypothesis.” Instead, our algorithms can tolerate label noise (alongside covariate noise) and achieve the information-theoretically optimal clean error under such a model (Proposition˜1.5).
@@
Learning with heavy contamination can be thought of as a search version of testable learning. More specifically, in testable learning the goal is to decide whether the input dataset is structured enough so that a near-optimal hypothesis can be found efficiently, while in HC-learning the goal is to find a subset of the input that is structured enough. An analogous connection was observed in [49] between TDS learning [78] and PQ learning [50] in the context of learning under distribution shift. There, the goal was to either decide whether the (unlabeled) test examples come from a distribution that is similar to the one the learner has trained on (TDS), or find a subset of the test examples where the learner is confident in its predictions (PQ). See Remark˜4.5.
@@ 2 Notation
We consider a $d$-dimensional feature space $\mathcal{X}$ which will either be ${\mathbb{R}}^{d}$ or the hypercube $\{\pm 1\}^{d}$. A polynomial $p$ over ${\mathbb{R}}^{d}$ is of the form $p(\mathbf{x})=\sum_{\alpha\subseteq{\mathbb{N}}^{d}}c_{p}(\alpha)\,\mathbf{x}^{\alpha}$, where $\mathbf{x}^{\alpha}=\prod_{i\in[d]}x_{i}^{\alpha_{i}}$ is a monomial of degree $\|\alpha\|_{1}$. The degree of $p$ is equal to the maximum $\|\alpha\|_{1}$ such that $c_{p}(\alpha)\neq 0$. Over $\{\pm 1\}^{d}$ we use the multilinear expansion $p(\mathbf{x})=\sum_{\mathcal{I}\subseteq[d]}c_{p}(\mathcal{I})\mathbf{x}^{\mathcal{I}}$, where $\mathbf{x}^{\mathcal{I}}=\prod_{i\in\mathcal{I}}x_{i}$. In both cases, we call $\mathbf{c}_{p}$ the vector of coefficients. We consider sets of points $S$ to contain examples that are separate instances of the correponding elements of $\mathcal{X}$. We denote with $\bar{S}$ the corresponding labeled set of examples. We use $\mathbf{x}\sim S$ to say that $\mathbf{x}$ is drawn uniformly from $S$. For a labeled distribution $\bar{\mathcal{D}}$ over $\mathcal{X}\times\{\pm 1\}$, $\mathcal{D}$ is the marginal on $\mathcal{X}$. In the following the distribution-specific algorithms have sample access to the unlabeled target distribution $\mathcal{D}^{*}$. We use $\mathcal{N}_{d}=\mathcal{N}(0,\mathbf{I}_{d})$ to denote the standard $d$-dimensional Gaussian and $\mathrm{Unif}_{d}=\mathrm{Unif}\{\pm 1\}^{d}$ for the uniform over the $d$-dimensional hypercube.
@@ 3 Iterative Polynomial Filtering
We say that a distribution $\mathcal{D}$ over $\mathcal{X}$ is A{A}-hypercontractive with respect to polynomials for some ${A}\geq 1$ if for any polynomial $p$ over $\mathcal{X}$ and any $t\geq 2$ we have
@@
Let $\mathcal{D}^{*}$ be a A{A}-hypercontractive distribution over a $d$-dimensional space $\mathcal{X}$. Consider parameters $\epsilon,\delta\in(0,1)$ and $R,\ell,m\geq 1$, and let $S_{\mathrm{inp}}$ be an arbitrary set of examples in $\mathcal{X}$ and $S_{\mathrm{ref}}$ a set of $m_{\mathrm{ref}}$ i.i.d. examples from $\mathcal{D}^{*}$. For a sufficiently large universal constant $C\geq 1$, if $m_{\mathrm{ref}}\geq\frac{R^{2}(CAd)^{2\ell}}{\epsilon^{3}}(\log\frac{1}{\delta})^{4\ell+1}$, then Algorithm˜1 on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m,\ell,R,\epsilon)$ runs in time $\operatorname{poly}(|S_{\mathrm{inp}}|,m_{\mathrm{ref}},(d+1)^{\ell})$ and outputs $S_{\mathrm{filt}}\subseteq S_{\mathrm{inp}}$ such that the following hold:
@@
1
21exLet $\beta\leftarrow 2(2A)^{2\ell}$; $\gamma\leftarrow\frac{\epsilon}{2R}$; $B\leftarrow 4(d+1)^{\frac{\ell}{2}}(\frac{\beta}{\epsilon})^{\frac{1}{2}}$; $\Delta\leftarrow\frac{\epsilon}{2B}$;
@@
We now give the proof of our main technical tool, Theorem˜3.2. The proof follows the approach of [76], but has a number of technical differences. In particular, the algorithm of [76] only gives guarantees for non-negative polynomials and works with respect to the uniform distribution on the hypercube. Here, we preserve the expectation of any polynomial, with respect to any target hypercontractive distribution that can be sampled efficiently.
Proof of Theorem˜3.2.
We first observe that the family $\mathcal{P}$ of polynomials $p$ of degree at most $\ell$ for which $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p(\mathbf{x})|]\leq\gamma$ and $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[(p(\mathbf{x}))^{2}]\leq\beta$ can be described by $O(m_{\mathrm{ref}})$ linear constraints plus one convex quadratic constraint over the coefficient vectors whose dimension is $(d+1)^{\ell}$. Therefore, lines 1 and 6 can be implemented as convex programs in time $\operatorname{poly}(|S_{\mathrm{inp}}|,m_{\mathrm{ref}},(d+1)^{\ell})$. Lines 1 and 1 can be implemented with a single pass of the input points.
@@
$\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[|p^{*}(\mathbf{x})|]$ $=\int_{\tau=0}^{\infty}\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}[|p^{*}(\mathbf{x})|>\tau]\;d\tau=\int_{\tau=0}^{B}\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}[|p^{*}(\mathbf{x})|>\tau]\;d\tau$ (3.1)
$\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p^{*}(\mathbf{x})|]$ $=\int_{\tau=0}^{\infty}\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p^{*}(\mathbf{x})|>\tau]\;d\tau\geq\int_{\tau=0}^{B}\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S_{\mathrm{ref}}}[|p^{*}(\mathbf{x})|>\tau]\;d\tau$ (3.2)
@@
We will show that with high probability over the clean and reference datasets, the number of removed points from $S_{\mathrm{cln}}$ is small. First, we account for the initial filtering of line 1. In this step, we remove points $\mathbf{x}$ from $S_{\mathrm{inp}}$ that give large absolute values to polynomials in $\mathcal{P}$, i.e., polynomials who, in particular, have bounded second norms over $S_{\mathrm{ref}}$. Due to Lemma˜A.2 and since $m_{\mathrm{ref}}\geq(CAd)^{2\ell}(\log\frac{1}{\delta})^{4\ell+1}$, the probability that some $\mathbf{x}$ drawn from $\mathcal{D}$ gives $|p(\mathbf{x})|>B$ for some $p\in\mathcal{P}$ is at most $\epsilon/4$. Therefore, the total number of points removed from $S_{\mathrm{cln}}$ in this step follows the binomial distribution with $m$ number of trials and probability of success at most $\epsilon/4$. By a standard Chernoff bound, and since $m\geq\frac{C}{\epsilon}\log(1/\delta)$, with probability at least $1-\delta/2$ we have
@@
Since $m,m_{\mathrm{ref}}\geq C^{\prime}\frac{(d+1)^{\ell}+\log(1/\delta)}{(\Delta/R)^{2}}$ for some sufficiently large universal constant $C^{\prime}\geq 1$ and the VC dimension of the class of polynomial threshold functions of degree $\ell$ is at most $(d+1)^{\ell}$, we have that with probability at least $1-\delta/2$ over $S_{\mathrm{cln}}$, $S_{\mathrm{ref}}$ the following holds for all polynomials $p$ of degree at most $\ell$ and all $\tau\geq 0$
@@
since $S_{\mathrm{cln}}$ and $S_{\mathrm{ref}}$ are both i.i.d. samples from the distribution $\mathcal{D}^{*}$ and $\operatorname{\mathbbm{1}}\{|p(\mathbf{x})|>\tau\}$ is equal to the sum of two degree-$\ell$ polynomial threshold functions, $\operatorname{\mathbbm{1}}\{p(\mathbf{x})>\tau\}+\operatorname{\mathbbm{1}}\{p(\mathbf{x})<-\tau\}$. Therefore:
@@
So far, we have proven the first part of Theorem˜3.2. We will now prove the second part, by showing the following claim.